Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] DURC_{Mark Dennehy} 10m Air Rifle Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Round 6 Due Date : [October 5]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DURC Mark Dennehy 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC_{Mark Dennehy} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30] DURC_{Mark Dennehy} 10m Air Rifle

Round 1 Due Date : [July 27]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Steven Watterson

10m Air Rifle Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Round 6 Due Date : [October 5]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Steven Watterson

10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Steven Watterson

10m Air Rifle

Round 1 Due Date : [July 27]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] DURC_{Ben Jones} 10m Air Rifle Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Round 6 Due Date : [October 5]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DURC Ben Jones 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC_{Ben Jones} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30] DURC_{Ben Jones} 10m Air Rifle

Round 1 Due Date : [July 27]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DRC_{J.K.Walsh} 10m Air Rifle

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Round 6 Due Date : [October 5]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]  DRC J.K.Walsh 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DRC_{J.K.Walsh} 10m Air Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DRC_{J.K.Walsh} 10m Air Rifle